Light of wavelength 560 nm goes through a pinhole of diameter 0.20 mm and falls on a wall at a distance of 2.00 m. What will be the radius of the central bright spot formed on the wall?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. For the thin oil film, d = 1 * 10^–4 cm = 10^–6 m, μ base oil = 1.25 and μ base x = 1.50 λ = 2μd/(n + 1/2) 2 * 10^-6 * 1.25 * 2/2n + 1 = 5 * 10^-6m/2n + 1 ⇒ λ = 5000 nm/2n + 1 For the wavelengths in the region (400 nm – 750 nm) When, n = 3, λ = 5000/2 * 3 + 1 = 5000/7 = 714.3 nm When, n = 4, λ = 5000/2 * 4 + 1 = 5000/9 = 555.6 nm When, n = 5, λ = 5000/2 * 5 + 1 = 5000/11 = 454.5 nm